Based on a survey, the mean weekly allowance of the Grade 11 students of XYZ

University is Php750.00 with a standard deviation of Php100.00. Assuming that

the weekly allowance is normally distributed, what is the z-score of a student with

a weekly allowance of:

a. Php950.00

b. Php600.00



1
Expert's answer
2021-11-17T00:38:46-0500

a.

"z=\dfrac{950-750}{100}=2"

b.


"z=\dfrac{600-750}{100}=-1.5"